OpenStudy (anonymous):

you and 4 friends go to a concert in how many different ways you can sit in the assigned seats? a.24 b.20 c.48 d.120

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Assuming there are only 4 seats for the 4 people then there are a total of \[4!=4 \times 3 \times 2 \times 1 = 24\] possible ways to sit. They way to think about this is: there are 4 choices for who can sit in the first seat, then 3 choices for the second seat and so on. When you get to the last seat there is only one person left and she has to sit there.

OpenStudy (anonymous):

so its not c

OpenStudy (anonymous):

oh k i got it wrong for the first time it 24